# Quantum-centric Supercomputing for Materials Science: A Perspective on Challenges and Future Directions

### v1

@@ II.6.2 Circuit optimization
II.7 Real-time classical processing of quantum information (within coherence time)
In this model of hybrid computation, the classical part is embedded or coupled with the quantum program and runs simultaneously with the quantum operations within the coherence time. This can significantly reduce the data-movement between the classical and quantum processors while providing the desired flexibility and short timing in controlling the quantum state. With mid-circuit measurement [309, 310], classical computation can use previous quantum operation outcomes to profile the quantum execution status, predict and even adjust the quantum operations for future iterations or steps. Some examples are QEC [311], progressive projection filtering [312, 313] and random walk phase estimation [314]. The key point is that the classical computation has to be performed while the quantum state is still coherent. Given the short coherence time for near-term devices, constraints on the latency of the classical computation can be stringent, possibly demanding a strongly coupled classical hardware accelerator (e.g., FPGA, GPU, or ASIC) physically in place where classical control is happening.
II.8 Error mitigation
The output of quantum devices is inevitably affected by noise. While QEC can theoretically be used to remove error from any computation once error rates of the physical devices are pushed below the fault-tolerance threshold, in the near term various techniques can be used to reduce the effect of noise on computation, generally referred to as quantum error mitigation (QEM) [43, 315, 191, 316]. When a noisy quantum device is used to compute expectation values of quantum observables, this in general results in a biased estimator. The goal of QEM is to minimize the bias and variance in expectation values from several different runs of the noisy quantum circuit using classical post-processing.
@@ II.8 Error mitigation
An alternative approach to QEM is called zero-noise extrapolation (ZNE). ZNE is based on the collection of noisy expectation values for various amplified values of the noise parameter. The expectation value at the zero-noise level is then extrapolated using a linear, polynomial, or exponential fit. ZNE can be implemented using analog (relying on pulse stretching [318]) or digital amplification methods (such as sub-circuit repetition [319, 320, 321]) that do not require a precise characterization of the noise, or using probabilistic error amplification (PEA) [322], which achieves the desired noise factor by randomly drawing circuits from a distribution implementing a rescaled Pauli noise model learned from the device. ZNE does not suffer from the severe sampling overhead of PEC, but it does not produce a guaranteed unbiased estimator. Its success depends on a proper choice of the fitting method (methods for systematically choosing the most appropriate one have been proposed, e.g., [321]), amplification factors for the noise (e.g., too large amplification will destroy the signal, too small amplification will make the extrapolation less precise), as well as the amplification method itself. Recently, ZNE has also been extended to quantum error correction [323].
@@
III Workload management
III.1 Overview
@@
In quantum-centric supercomputing systems, computational workloads can also be defined based on Quantum-HPC integration types, such as HPC-for-Quantum, Quantum-about-HPC, Quantum-in-HPC scenarios, involving interplay between quantum and classical tasks. Consistency in Quantum-x-HPC or HPC-x-Quantum seems desirable and possible. Quantum-about-HPC workload involves quantum tasks that require HPC resources for pre- and post-processing. In this integration type, the quantum tasks are independent of the HPC tasks, and the HPC resources are used to support the quantum tasks. Quantum-in-HPC workload involves HPC tasks that require quantum resources for acceleration, where the quantum tasks are tightly coupled with the HPC tasks, and the HPC resources are used to support the quantum tasks.
Finally, HPC-for-Quantum workload involves tight integration with quantum resources in real-time, within the coherence time of the quantum system II.7. The classical resources are used to support the quantum tasks, and the classical tasks are tightly coupled with the quantum tasks [325]. Circuits with classical control often referred as dynamic circuits.
@@ III.2 Resource management
III.2.1 Resource model
A resource model refers to the framework or method used to manage and allocate computational resources. Specifically, in this section, we explore the distinction and relationship between Quantum Processing Units (QPUs) and qubits, the placement and location of quantum and classical resources and ownership models.
@@ III.2.1 Resource model
The introduction of quantum processing units (QPUs) marks a significant departure from the conventional approach to traditional high-performance computing (HPC) scaling. Instead of pursuing the conventional path of continually improving semiconductor processors for greater speed and density, or incorporating specialized processors like GPUs, and optimizing their coordinated use through parallelization, QPUs take a fundamentally different approach. While one might assume that QPUs merely serve as an additional component, they fundamentally alter the way data is represented, employing the language of quantum circuits governed by the principles of quantum mechanics. This constitutes a distinct computational model, diverging from all that has come before it. However, it’s essential to note that QPUs are not designed for general-purpose computing tasks. Consequently, practical implementations necessitate the careful orchestration of both classical and quantum computation resources.
@@
Co-located resources refer to computational resources physically situated within the same facility or data center as the user or workload management system. There are several advantages of having co-located quantum and classical resources. Minimal network latency and high-speed data transfer capabilities, which is especially advantageous for applications where access to large amounts of (low-level) data (like counts) is required. Full control and management of co-located resources, enabling customization and configuration based on specific requirements. Data security. Co-located resources location implements on-premises integration type (i) III.1.
Remote quantum resources, on the other hand, are located at a physical distance from HPC classical infrastructure, often in off-site data centers, cloud environments (cloud model integration type (ii) III.1), or partner organizations. Remote resources extend the reach and capabilities of HPC systems, but they come with unique characteristics. Elasticity is one of them, allowing for scaling up and down quantum resources based on need. Another characteristic is sharing. At this moment, it is a case when a quantum processor is shared among multiple users, which drives down the cost of execution, but also introduces variation in performance due to resource contention. Network overhead can impact data transfer rates and communication latency, especially in applications with large loads.
@@
High-performance computing facilities generally serve large numbers of users with varying computational resource requirements for their computational workloads (jobs) II. Users are part of user groups and projects, which have assigned queues. A project has some resource allocation quota (e.g., $10^{6}$ node hours), whereas queues are defined by criteria such as priority and the amount of computational resources available per job (e.g., $n$ CPU cores, $m$ GPUs) to a user of that queue. At the level of user groups and projects, the integration of quantum information processing devices with classical HPC computational resources is straightforward: access to a quantum resource could be granted to a given user assigned to a qpu_user group with project quota $Q$ specifying the number of “quantum node” hours.
@@ III.2.2 Job Scheduling
Common frameworks in use today within HPC facilities and cloud platforms include SLURM [328], TORQUE [329], Altair Grid Engine [330], and IBM Platform LSF [331], which provide highly scalable control over heterogeneous computer clusters and jobs submitted to them. To execute a workload, a user configures their job using a submission script based on the syntax of the associated scheduler. Submission scripts essentially define a job: project name, resource requirements (e.g., node count, CPU count, memory size, and wall clock limit), tasks to execute, input and output data paths, etc. Note the governance of an HPC system (see Sec. III.4) could require all input data be readily available, e.g., on a scratch area accessible only by the user. These scripts are submitted from a login (head) node to the WMS or job scheduler, where they are queued and finally executed on one or more compute nodes.
@@
Additionally, quantum-centric supercomputing workload scheduling must also consider the coupling between the quantum and classical tasks and assign and co-locate tasks to the resources accordingly. Different quantum-classical coupling types represent workloads with unique requirements and challenges, such as the lack of a unified standard for accessing QPUs and expressing hybrid workloads, and the need for parallelization of quantum workloads across nodes, cores, and accelerators. For example, classical tasks tightly coupled to quantum tasks, e.g., for the HPC-for-Quantum integration type, must be co-allocated and assigned to nearby QPUs and GPUs. Additionally, scheduling of QPU jobs must be prioritized to ensure the optimal utilization of the QPUs while minimizing the time-to-solution and energy-to-solution [325] (section III.2.2).
@@
Cloud queues. There is a practice nowadays of cloud vendors are providing access to quantum services that are executing quantum payload. In order to provide fair allocation of resources, cloud providers are implementing their own queuing mechanisms behind cloud endpoints. It is important to understand that cloud vendors have different job queuing and priority implementations, and all might employ a shared resource ownership model III.2.1.
@@
Endpoint
Request Body
Response
Description
GET allocations/
allocation_ids: list(str) fields: list(str)
allocations: list(dict)
Fetch resource allocations listed by id.
POST allocations/
specification: dict
allocation_id: str
This allocates a set of resources from a resource request specification.
DEL allocations/
allocation_id: str
status: bool
Destroy a resource allocation and return resources to the pool of available resources.
GET resources/
specification: dict fields: list(str)
resources: list(dict)
Fetch the current state of resources on a system. This should report status of anything that a user may be able to specify for their job placement, including QPU status and software stack status
GET tasks/
task_ids: list(str) fields: list(str)
tasks: list(dict)
If a list of tasks are passed in, provide information on currently running tasks on the system. If no tasks passed in, provide information on all running tasks. This may constrain the results for the executing user so that it shows only their own jobs if the user is otherwise unprivileged.
POST tasks/
task_information: dict allocation_id: str batch_id: str
allocation_id: str
Launch a task.
DEL tasks/
task_ids: list(str) signal: int
status: bool
Send a signal to tasks for termination.
@@
III.3 Middleware
Quantum computing is evolving towards modular architectures comprising multiple quantum processing units (QPUs) coupled to classical computing nodes. Moreover, emerging applications that aim to benefit from quantum acceleration involve significant classical and quantum computational components, such as different phases of data collection and streaming, parameter tuning, simulation, and analysis. Typically, such quantum and classical coupling depend upon interaction between application components within and outside the coherence time of the quantum system, i.e., on tight- and medium-coupling across multiple quantum computing units (QPUs) or loose-coupling in a workflow application [325] (see section III.1). Hence, there is a need to design middleware systems that can facilitate the efficient understanding and interplay between quantum and classical components in an end-to-end workflow. Regarding quantum-centric supercomputing systems, middleware should also leverage well-established high-performance computing abstractions and must be compatible with existing HPC software stacks for managing hybrid workloads, tasks, and resources to integrate quantum computing into HPC systems seamlessly.
@@ III.4.2 Policies
IV Classical simulation of quantum systems
IV.1 Why use classical simulators

### v2

@@ III.6.2 Circuit optimization
III.7 Real-time classical processing of quantum information (within coherence time)
In this model of hybrid computation, the classical part is embedded or coupled with the quantum program and runs simultaneously with the quantum operations within the coherence time. This can significantly reduce the data-movement between the classical and quantum processors while providing the desired flexibility and short timing in controlling the quantum state. With mid-circuit measurement [310, 311], classical computation can use previous quantum operation outcomes to profile the quantum execution status, predict and even adjust the quantum operations for future iterations or steps. Some examples are quantum error correction (QEC) [312], progressive projection filtering [313, 314] and random walk phase estimation [315]. The key point is that the classical computation has to be performed while the quantum state is still coherent. Given the short coherence time for near-term devices, constraints on the latency of the classical computation can be stringent, possibly demanding a strongly coupled classical hardware accelerator – e.g., field programmable gate array (FPGA), graphics processing unit (GPU), or application-specific integrated circuit (ASIC)– physically in place where classical control is happening.
III.8 Error mitigation
The output of quantum devices is inevitably altered by noise throughout a computation. While QEC can theoretically be used to remove error from any computation once error rates of the physical devices are pushed below the fault-tolerance threshold, in the near term various techniques can be used to reduce the effect of noise on computation, generally referred to as quantum error mitigation (QEM) [43, 316, 192, 317]. When a noisy quantum device is used to compute expectation values of quantum observables, this in general results in a biased estimator. The goal of QEM is to minimize the bias and variance in expectation values from several different runs of the noisy quantum circuit using classical post-processing.
@@ III.8 Error mitigation
An alternative approach to QEM is called Zero-Noise Extrapolation (ZNE). ZNE is based on the collection of noisy expectation values for various amplified values of the noise parameter. The expectation value at the zero-noise level is then extrapolated using a linear, polynomial, or exponential fit. ZNE can be implemented using analog (relying on pulse stretching [319]) or digital amplification methods (such as sub-circuit repetition [320, 321, 322]) that do not require a precise characterization of the noise, or using probabilistic error amplification (PEA) [323], which achieves the desired noise factor by randomly drawing circuits from a distribution implementing a rescaled Pauli noise model learned from the device. ZNE alleviates sampling overhead of PEC, however does not guarantee an unbiased estimator. Its success depends on a proper choice of the fitting method (methods for systematically choosing the most appropriate one have been proposed, e.g., [322]), amplification factors for the noise (e.g., too large amplification will destroy the signal, too small amplification will make the extrapolation less precise), as well as the amplification method itself. Recently, ZNE has also been extended to quantum error correction [324].
@@
IV Workload management
IV.1 Overview
@@
In QCSC systems, computational workloads can also be defined based on Quantum-HPC integration types, such as HPC-for-Quantum, Quantum-about-HPC, Quantum-in-HPC scenarios, involving interplay between quantum and classical tasks. Consistency in Quantum-x-HPC or HPC-x-Quantum seems desirable and possible. Quantum-about-HPC workload involves quantum tasks that require HPC resources for pre- and post-processing. In this integration type, the quantum tasks are independent of the HPC tasks, and the HPC resources are used to support the quantum tasks. Quantum-in-HPC workload involves HPC tasks that require quantum resources for acceleration, where the quantum tasks are tightly coupled with the HPC tasks, and the HPC resources are used to support the quantum tasks.
Finally, HPC-for-Quantum workload involves tight integration with quantum resources in real-time, within the coherence time of the quantum system III.7. The classical resources are used to support the quantum tasks, and the classical tasks are tightly coupled with the quantum tasks [326]. Circuits with classical control often referred as dynamic circuits.
@@ IV.2 Resource management
IV.2.1 Resource model
A resource model refers to the framework or method used to manage and allocate computational resources. Specifically, in this section, we explore the distinction and relationship between QPUs and qubits, the placement and location of quantum and classical resources and ownership models.
@@ IV.2.1 Resource model
The introduction of QPUs marks a significant departure from the conventional approach to traditional HPC scaling. Instead of pursuing the conventional path of continually improving semiconductor processors for greater speed and density, or incorporating specialized processors like GPUs, and optimizing their coordinated use through parallelization, QPUs take a fundamentally different approach. While one might assume that QPUs merely serve as an additional component, they fundamentally alter the way data is represented, employing the language of quantum circuits governed by the principles of quantum mechanics. This constitutes a distinct computational model, diverging from all that has come before it. However, it’s essential to note that QPUs are not designed for general-purpose computing tasks. Consequently, practical implementations necessitate the careful orchestration of both classical and quantum computation resources.
@@
Co-located resources refer to computational resources physically situated within the same facility or data center as the user or workload management system. There are several advantages of having co-located quantum and classical resources. Minimal network latency and high-speed data transfer capabilities, which is especially advantageous for applications where access to large amounts of (low-level) data (like counts) is required. Full control and management of co-located resources, enabling customization and configuration based on specific requirements. Data security. Co-located resources location implements on-premises integration type (i) IV.1.
Remote quantum resources, on the other hand, are located at a physical distance from HPC classical infrastructure, often in off-site data centers, cloud environments (cloud model integration type (ii) IV.1), or partner organizations. Remote resources extend the reach and capabilities of HPC systems, but they come with unique characteristics. Elasticity is one of them, allowing for scaling up and down quantum resources based on need. Another characteristic is sharing. At this moment, it is a case when a quantum processor is shared among multiple users, which drives down the cost of execution, but also introduces variation in performance due to resource contention. Network overhead can impact data transfer rates and communication latency, especially in applications with large loads.
@@
High-performance computing facilities generally serve large numbers of users with varying computational resource requirements for their computational workloads (jobs) III. Users are part of user groups and projects, which have assigned queues. A project has some resource allocation quota (e.g., $10^{6}$ node hours), whereas queues are defined by criteria such as priority and the amount of computational resources available per job (e.g., $n$ CPU cores, $m$ GPUs) to a user of that queue. At the level of user groups and projects, the integration of quantum information processing devices with classical HPC computational resources is straightforward: access to a quantum resource could be granted to a given user assigned to a qpu_user group with project quota $Q$ specifying the number of “quantum node” hours.
@@ IV.2.2 Job Scheduling
Common frameworks in use today within HPC facilities and cloud platforms include SLURM [329], TORQUE [330], Altair Grid Engine [331], and IBM Platform LSF [332], which provide highly scalable control over heterogeneous computer clusters and jobs submitted to them. To execute a workload, a user configures their job using a submission script based on the syntax of the associated scheduler. Submission scripts essentially define a job: project name, resource requirements (e.g., node count, CPU count, memory size, and wall clock limit), tasks to execute, input and output data paths, etc. Note the governance of an HPC system (see Sec. IV.4) could require all input data be readily available, e.g., on a scratch area accessible only by the user. These scripts are submitted from a login (head) node to the WMS or job scheduler, where they are queued and finally executed on one or more compute nodes.
@@
Additionally, quantum-centric supercomputing workload scheduling must also consider the coupling between the quantum and classical tasks and assign and co-locate tasks to the resources accordingly. Different quantum-classical coupling types represent workloads with unique requirements and challenges, such as the lack of a unified standard for accessing QPUs and expressing hybrid workloads, and the need for parallelization of quantum workloads across nodes, cores, and accelerators. For example, classical tasks tightly coupled to quantum tasks, e.g., for the HPC-for-Quantum integration type, must be co-allocated and assigned to nearby QPUs and GPUs. Additionally, scheduling of QPU jobs must be prioritized to ensure the optimal utilization of the QPUs while minimizing the time-to-solution and energy-to-solution [326] (section IV.2.2).
@@
Cloud queues. There is a practice nowadays of cloud vendors are providing access to quantum services that are executing quantum payload. In order to provide fair allocation of resources, cloud providers are implementing their own queuing mechanisms behind cloud endpoints. It is important to understand that cloud vendors have different job queuing and priority implementations, and all might employ a shared resource ownership model IV.2.1.
@@
Endpoint Request Body Response Description
GET allocations/ allocation_ids: list(str) fields: list(str) allocations: list(dict) Fetch resource allocations listed by id.
POST allocations/ specification: dict allocation_id: str This allocates a set of resources from a resource request specification.
DEL allocations/ allocation_id: str status: bool Destroy a resource allocation and return resources to the pool of available resources.
GET resources/ specification: dict fields: list(str) resources: list(dict) Fetch the current state of resources on a system. This should report status of anything that a user may be able to specify for their job placement, including QPU status and software stack status
GET tasks/ task_ids: list(str) fields: list(str) tasks: list(dict) If a list of tasks are passed in, provide information on currently running tasks on the system. If no tasks passed in, provide information on all running tasks. This may constrain the results for the executing user so that it shows only their own jobs if the user is otherwise unprivileged.
POST tasks/ task_information: dict allocation_id: str batch_id: str allocation_id: str Launch a task.
DEL tasks/ task_ids: list(str) signal: int status: bool Send a signal to tasks for termination.
@@
IV.3 Middleware
Quantum computing is evolving towards modular architectures comprising multiple QPUs coupled to classical computing nodes. Moreover, emerging applications that aim to benefit from quantum acceleration involve significant classical and quantum computational components, such as different phases of data collection and streaming, parameter tuning, simulation, and analysis. Typically, such quantum and classical coupling depend upon interaction between application components within and outside the coherence time of the quantum system, i.e., on tight- and medium-coupling across multiple quantum computing units (QPUs) or loose-coupling in a workflow application [326] (see section IV.1). Hence, there is a need to design middleware systems that can facilitate the efficient understanding and interplay between quantum and classical components in an end-to-end workflow. Regarding quantum-centric supercomputing systems, middleware should also leverage well-established high-performance computing abstractions and must be compatible with existing HPC software stacks for managing hybrid workloads, tasks, and resources to integrate quantum computing into HPC systems seamlessly.
@@ IV.4.2 Policies
V Programming models
V.1 Introduction
A programming model describes and provides the mechanisms for instructing how a computer system is to perform a computation, based on the execution models of specific computing hardware and programming language. These can usually be separated because any underlying execution model for a given hardware platform may be made transparent in any given language through calls to functions through abstractions. For the purposes of this paper, we will adopt this common viewpoint and focus on language-independent programming models supporting developers in several roles working across a layered execution model. Likewise, we will assume independence of programming paradigm, allowing for its choice to be driven by the details of an individual project and its requirements.
A programming model for a heterogeneous computing system combining both classical and quantum elements can be described in terms of four layers, illustrated in Fig. 8, comprising an application’s execution and categories of developers who write code describing computational procedures across these layers.
Working up from the bottom, Layer-1, which we call the Hardware layer, is where hardware-dependent operations occur on specific devices. On the quantum side, the programming model at Layer-1 provides access to processors and other electronics, software, and firmware which directly control the quantum device, and on the classical side it provides processor-specific instructions according to its Instruction Set Architecture (ISA), itself an abstraction of the processor hardware.
Layer-2, which we call the Computation layer, is where low level computational flows and information are described. For quantum this means constructing and optimizing circuits which are then mapped onto specific quantum hardware by compilation (and transpilation), operating on quantum information and interfaces to classical information. On the classical side this means functions optimized for specific underlying processors including CPUs (of varying ISAs), GPUs via a manufacturer’s API, or other environments for specialized processors such as inference engines, processors aimed at AI problems, or custom FPGAs.
The Library or Function layer, Layer-3, is where computational capabilities in Layer-2 are abstracted into functions aimed at particular problem domains and the hardware-independent computations supporting them. These are used by application developers without necessarily being concerned with the details of the underlying execution model.
Finally, Layer-4, which we call the Application layer, is where domain-specific application developers construct programs and workflows that perform computations to solve problems relevant to domain experts, expressed in domain-relevant terms, without necessarily being concerned with the detailed mechanisms of those computations.
Figure 8: Programming model layers. (L1) Hardware layer: electronics, firmware V.5. (L2) Computational layer: circuits V.3 V.6. (L3) Library layer: functions. (L4) Application layer: domain specific applications.
V.2 Related Work
It’s useful and instructive to consider popular existing programming models and how they relate to this layering. Apache Spark [351] is a programming model for large-scale data processing whose base programming language is Java; SYCL [352], CUDA [353], and High Performance ParalleX (HPX) [354] are programming models developed in idiomatic C/C++, spanning Layer-2 and Layer-3.
There do exist domain-specific programming libraries such as OpenFermion [355], which facilitate the mapping of electron-structure problems to quantum circuits. While certainly useful, it is arguable whether this approach represents a complete programming model per se rather than a suite of helper routines for a specific workflow (module specification $\to$ integral generation $\to$ circuit mapping) at Layer-2.
SimuQ [356] defines a domain-specific language for modeling arbitrary Hamiltonians, compiles these to an abstract analog instruction set, then lowers that to a sequence of pulses rather than to a quantum circuit at Layer-2. Programming-model support for the direct expression of Hamiltonians may be beneficial in Layer-3 in defining functions for building materials-science applications. One challenge, however, is performance scalability of the software stack. Peng et al. [356] report SimuQ compilation times that grow rapidly with the number of sites in the Hamiltonian as well as a number of Hamiltonians that fail to map to the underlying quantum computer due to hardware constraints.
Message Passing Interface (MPI) [357] and Open Multi-Processing (OpenMP) [358] are open-standard specifications that aim generally for the programming of parallelized computations on heterogeneous classical computing platforms, and are examples operations spanning Layers 2 and 3.
Programming models can be directives-based, decorated, or provide extensions to their associated programming languages. In each case, a compiler capable of translating code written using a given programming model is needed to target specific backend hardware at Layer-1. A programming model incorporating semantics and execution model for quantum computing systems will naturally extend existing models targeting classical hardware. Importantly, a quantum computing-enabled programming model should provide abstractions such that applications can be developed independently from the underlying computing model or hardware; that is, at Layer-4 they should be agnostic with respect to both classical and quantum hardware. In this way, application developers can write software that can be executed on arbitrary hardware using functions in libraries at Layer-3, with possible customizations in Layer-2 if needed, and a backend compiler to translate the higher-level code (including language extensions, etc.) to the target machine code.
An initial prototype of such a programming model for the development of quantum computing applications is XACC/qcor [359, 360], an extensible language and classical-quantum compilation framework for Rigetti, IBM, and D-Wave quantum systems. XACC follows the co-processor model, akin to SYCL and CUDA, treating the quantum system as an computational accelerator, similar how a general purpose graphics processing unit (GPGPU) computing is used. This work paved the way for NVIDIA’s open-source CUDA Quantum platform [361] which extends their existing CUDA framework, providing support for programming QPUs in addition to CPUs and CUDA GPUs.
V.3 Execution Model
An execution model describes the behavior of an application and is exposed via a programming model. The hardware, or device, architectures help to define the execution models based on their capabilities and functionalities, and the programming model presents an abstract implementation of the underlying capabilities to enable general (or expected) purpose computation. In classical computing, a well known classification of architectures includes those found in Flynn’s taxonomy [362]. In this section, we complement Flynn in the context of quantum computation and how this classification can be incorporated into a programming model for hybrid classical-quantum applications.
Flynn’s taxonomy is laid out in two dimensions: the number of instruction streams or programs and the number of data streams. The single instruction, single data (SISD) model represents classical sequential machines following the von Neumann architecture. In the immediate term of quantum computing, until QPU partitioning or parallel QPU processing is possible, we foresee that SISD will be the predominant architecture. Programming languages exposing this execution model include C, C++, and Fortran, which are dominant languages within HPC. Parallel quantum computations, wherein, e.g., a large QPU can be partitioned into topologically equivalent subprocessors, single instruction, multiple data (SIMD) will be beneficial; SIMD capabilities can then be exploited to parallelize quantum computations. A recent popular addition to the SIMD family is GPUs, which manipulate multiple data through the same set of instructions executed across many processors in lock-step, known as single instruction, multiple threads (SIMT) [363].
Multiple instruction, multiple data (MIMD) is generally accepted to represent parallelism. Again, supporting an architecture within a programming model depends on the underlying hardware capabilities which are exposed to the user. Modern languages such as Erlang [364], Go [365], or Python [366], and recent revisions of C++ [367] present built-in architecture-dependent concurrency models, and parallelism is most often exploited either via extensions to traditional HPC programming models such as CUDA [368], OpenMP [369], OpenMPI [370], or via careful allocation of resources and explicit control flow declaration known as “workload orchestration”. Ray [371] is an example of a parallel workload orchestrator.
In our layered vision of programming models, at Layer-1 (i.e., Hardware Layer), classical hardware shows instances of several kinds of execution models: instruction-level parallel (superscalar) processors, instances of MIMD and GPUs, address SIMD and SIMT; Tensor Processing Units, and Data Processing Units support unloading specific workloads from CPUs, ASICS are examples of programmable technologies for specialized SISD; and novel applications can be explored through the use of reconfigurable hardware such as FPGAs. On the quantum side, general-purpose quantum computers can essentially be considered SIMD machines with a computational memory in which each qubit has both independent and combined processing capabilities.
At Layer-2 (Computation Layer), the hardware architecture is abstracted into the programming model’s execution model. This, in turn, will be reflected by the programming language subject to compilation. Language-specific statements control the underlying hardware through a human-friendly syntax and developer ergonomics, presenting an abstract SISD model. It is through the careful coordination that access to parallelism and other execution models is granted to higher layers. In quantum computing, although there exist programming languages capturing quantum semantics (e.g., OpenQASM 3 [372], Q# [373], and Quipper[374]), the control flow of quantum circuits, circa dynamic circuits, is simple enough to be captured as data and to be translated into hardware instructions mostly directly.
Layer-3 (Library Layer) takes advantage of Layer-2’s execution model libraries to access actual programming models beyond SISD variants. This layer increases the level of abstraction into libraries, frameworks, and software development kits (SDK), introducing concepts agnostic to the hardware execution model, and tools for interfacing with the abstract execution model. A classical example, in AI engineering, TensorFlow [375] exploits the concept of a graph, and implements a variant of the data-flow execution model, where operations are executed as soon as the input operands are ready. Tools such as TensorBoard [376] allow for the inspection and debugging of graphs. Similarly, on the quantum side, the Qiskit SDK [377] introduces concepts from quantum computing such as quantum registers, gates, and measurements, and captures the circuit execution model as a directed acyclic graph (DAG). The circuit and DAG visualizers [378] allow for the exploration of these quantum program representations.
Finally, Layer-4 makes use of these libraries, frameworks, and SDKs to implement domain-specific applications. Applications at this level act as coordinators, directing the workloads needed to solve specific problems. Within the quantum computing ecosystem, Quantum Serverless [291] can coordinate an execution workload through many clouds and quantum computing providers and thus, understand concepts coming from both the classical and quantum worlds. At this uppermost level, developers make use of the most suitable programming model based on trade-offs between speed, solution quality, and resource cost.
V.4 Quantum Kernels
In the context of classical computing, invocation of accelerator kernels can happen in several ways: (1) host (CPU) code; (2) strictly via other kernels (device-side, kernel launch); or (3) host or kernel code. Scheduling of host code and accelerator kernels similarly can be performed in three ways: (1) implicitly, if a programming model provides runtime mechanisms for managing a task DAG; (2) asynchronously, through non-blocking calls that return immediately after being called; or (3) explicitly, via synchronization points.
A programming model should enable quantum kernels (circuits, possibly with classical control flow) to execute in parallel, synchronously and asynchronously with respect to the kernels themselves and classical computations in order to cover a wide variety of use cases. This type of execution model is akin to that of classical accelerator (e.g., GPU) programming wherein asynchronous accelerated kernels can be desirable but data dependencies permit only synchronous execution. However, a major difference between purely classical and classical-quantum hybrid execution are the latencies involved; for instance, classical accelerators interface via high-speed buses whereas today’s free and publicly available quantum devices are generally cloud-based.
The rate at which quantum computations can be performed (i.e., the analog to classical CPU clock speed) is set by the qubit technology which range from $O(1)$ Hz (photonic qubits) to $O(10^{6})$ Hz (superconducting qubits). Programming models must provide kernel scheduling mechanisms, or runtime, for orchestrating devices where latencies can range from $O(10^{-6})$ seconds to $O(1)$ minutes or greater. In cases where classical and quantum computations are decoupled — i.e., independent classical and quantum information processing with no communication between processes is needed — no orchestration would be required from the programming model. Indeed, classical and quantum routines or applications can be developed in separate languages and executed independently. In this section, we are concerned with ‘near-time’ hybrid computations wherein classical and quantum resources work in unison to complete a task.
Scheduling and execution of quantum kernels can be handled by a queuing system within the programming model. Similar to the SYCL programming model [352], a queue could have a one-to-one correspondence with a device; multiple queues would be necessary to work with multiple devices. Once a queue has associated to it a quantum device, that device can then be queried directly from the queue to obtain device-specific information, e.g., number of qubits and supported features. A quantum extension to the queue model would include sampling functionality to specify the number of executions (or shots) of a given kernel. Furthermore, a queue would capture the necessary semantics for execution, data input and output, etc., and provide within its API synchronous and asynchronous capabilities.
In principle, quantum accelerator kernels should be indistinguishable from classical accelerator kernels. That is, the semantics should be such that both classical and quantum kernel can expressed in the same way. For example, classical kernels generally define an “index space” over which an instance of the kernel executes. Each kernel instance executes the same code but may operate on different data or traverse a different execution pathway through its code. This mechanism, coupled to embarrassingly parallel workloads, is what empowers GPUs in today’s HPCs. For a quantum device kernel, the concept of an index space can be extended to include device (qubit) partitioning wherein, e.g., a single circuit instance can be mapped multiple times to the topology of a large device and hence execute in parallel.
In addition to partitioning, various quantum “backends” (i.e., control electronics, classical computational devices, etc.) will likely have different technical capabilities; e.g., control-flow, dynamical decoupling schemes, etc. Such capabilities could be configurable so a developer may choose the schemes most appropriate for their kernel. Therefore, an API specification should permit extensibility for vendors to expose all desired functionality for their backend devices via a common plugin mechanism, allowing for interoperability between different vendor library implementations.
In HPC, vendor libraries are typically written in low-level systems languages (Layer-2) such as C and C++ which are the de facto for applications targeting a set of heterogeneous architectures on current HPC platforms. A programming model built atop these would require language extensions for allow the expressiveness necessary to write circuit applications. Language extensions also allow for a single-source programming model such that development would be streamlined and eliminate the interplay of multiple languages (and thus compilers, interpreters, etc.). Moreover, this makes classical control-flow within quantum kernels more natural for developers. An alternative to a single large-scale application is an admixture of distinct executables whose orchestration is handled by a workflow management system (WFMS), see  [379]. The WFMS is effectively a wrapper around multiple executables — that may or may not have intercommunication — and steers their execution based on a predefined DAG or logic that enables dynamism of the overall flow.
V.5 Memory Model
Quantum processors operate at a very particular level in terms of memory and, in fact, a quantum processor can be regarded as a computational memory. In contrast with typical classical formalism where operations usually happen outside dedicated memory, quantum computing operates on the data input and the lifetime of quantum information is limited to the duration of the circuit carrying out the computation, with some notable exceptions like inherently quantum memories (QRAM and QROM). These quantum memories, hitherto never realized, would naturally span layers 1 and 2, and could play a fundamental role in classical data loading for some algorithms in which the implementation of the quantum oracle is critical to retain any level of quantum advantage [380]. Section LABEL:classical-data-loading provides a thorough discussion on this subject.
On a more classical standpoint, and similar to the concept of shared memory in HPC architectures, one could envision a model whereby some memory is shared at the QPU controller level with other processing units (GPUs, CPUs) in order to exploit different aspects of the quantum pre- and post-processing tasks, as well as fast loading of quantum circuits or parts of circuits.
V.6 Data Types
The integration of classical computing resources with QPUs involves necessarily some common interface for data consumption that bridges largely different DSLs. A seamless workflow that exploits both CPUs/GPUs/xPUs and QPUs will include steps to transform data types as needed by each unit. It is worthwhile to differentiate between data interpreted as instructions and other type of data like metadata or data carrying computing information. Within the scope of programming models, the quantum nature of the information existing in the QPU during the computation is unimportant. Any interaction of the QPU with supporting or co-computing units before, during, or after the quantum computation occurs largely in the form of classical information. In fact, both the inputs and outputs for quantum processors are naturally classical except for a few notable exceptions (QRAM being one example, as discussed in Section III.5). The natural outputs from QPUs are probability distributions, to which we have a level of access limited by our ability to sample them. Certain quantum algorithms like quantum phase estimation or Shor’s algorithm produce distributions so peaked that one or only a few samples are enough to provide the answer. Other algorithmic approaches based on heuristic methods, however, produce output distributions from which the value of certain quantum observable is computed by repeated sampling. This value can either represent the answer to the quantum subroutine, or it can be passed back to other -classical- computing resources to generate a new input to the QPU.
Inputs to quantum processors are, at the highest level of abstraction (Layer-2), a series of operations, that can be both unitary and non-unitary, and which can be represented by a quantum circuit. These quantum circuits can then take many levels of specificity throughout the transpilation and compilation chains, from the multilevel representation that a quantum assembly language like OpenQASM provides to the binaries used by the classical control hardware to implement the quantum gates on the quantum processor. These gates, or instructions, are parameterized to get instanced by data from the algorithm input. The gates are ultimately expressed as physical pulses of energy that interact with the quantum hardware (Layer-1). Thus, these pulses ultimately contain both instruction data and computational data. It is at higher levels of abstraction, however, where HPC could possibly help with circuit optimization, transpilation, and routing, working with circuit representations involving data types digestible by the DSL of the different processing units.
VI Classical simulation of quantum systems
VI.1 Why use classical simulators